Find the equation of a line that is perpendicular to the line y=1/9x+3 ad contains the point (-3,0)
Answer by lwsshak3(11628)

Find the equation of a line that is perpendicular to the line y=1/9x+3 ad contains the point (-3,0)
**
equation for a straight line:y=mx+b, m=slope, b=y-intercept
slope of given line=1/9
slope of line perpendicular to given line=-9 (negative reciprocal)
equation: y=-9x+b
solve for b using coordinates of given point (-3,0)
0=-9*-3+b
b=-27
equation: y=-9x-27